OpenStudy (caitlinnr14):

Solve: −4(x + 10) − 6 = −3(x − 2) −40 −46 −52 52

OpenStudy (uri):

−4(x + 10) − 6 = −3(x − 2) -4x-40 -6=-3x+6 -4x+3x=6+40+6 -x=52 =-52

OpenStudy (caitlinnr14):

wow thx @uri

OpenStudy (uri):

Yw :)